# Can AI writing be salvaged? Mitigating Idiosyncrasies and Improving Human-AI Alignment in the Writing Process through Edits

### v3

@@ 4.5. Collecting Edits on LLM Generated responses
With the finalized taxonomy of edits, we next conducted a larger-scale annotation study. The purpose of this study was to collect edits from writers on LLM-generated responses, categorizing them according to the established taxonomy.
@@
After completing their edits, participants assigned two scores to the sample: an Initial Writing Quality Score (IWQS) for the original response quality, and a Final Writing Quality Score (FWQS) for the post-edit quality. Both used a 1-10 scale, with 1 being lowest and 10 being the highest quality. The scores were incorporated to add a quantitative dimension to the qualitative process of editing. Additionally, the self-reported writing quality scoring system serves as a signal for writers to recognize their own improvements, set personal goals, and develop intrinsic motivation for enhancing their work.
@@ 5.1. Overall Statistics
Figure 5. (5(a)) the categories of edits they implement, and (5(b)) the relationship between writing quality scores and error categories.
We analyze the editing process by examining edit operations: insertion, deletion, or replacement. An edit is an insertion if it deletes no characters and adds 40+ characters net. Conversely, it’s a deletion if it adds no characters and removes 40+ characters net. All other edits are replacements. Figure 4(a) shows edit operations by participant for each paragraph. Replacements are most frequent (74%), followed by deletions (18%) and insertions (8%). Editing styles vary: some participants primarily use replacements (W2, W9, W10, W16, W18), while others employ deletions more often (W1, W5, W7, W8, W13, W17). Insertions are uncommon across all participants. To quantify meaning-preserving vs. meaning-changing edits, we calculate semantic similarity between original and edited text using BERT score(Zhang et al., 2019). Using a threshold of 0.6  [^7] This threshold was decided by manually analyzing 100 edits, we classify edits with similarity ¿ 0.6 as meaning-preserving. Of 6468 non-deletion edits, 70% are meaning-preserving, with the rest meaning-changing. This finding supports our Design Principle 2.
@@ 5.2. Writers differ greatly in the amount of editing they do: But to what extent?
Table 7 shows how sometimes writers select the exact same problematic span but assign different categories. For a span both W3 and W16 selected “,the numbers glaring back at me like an unsolvable riddle” (Table 7 Row 1) but chose separate categories (Table 7 Row 2). Similarly, W3 and W16 both choose the exact same span, but different categorizations. It should however be noted that both categorizations can be correct interpretations. When one relies on overused phrases or clichés, they often state the obvious or provide information that readers can easily infer implicitly. This results in redundant or superfluous exposition that doesn’t add value to the narrative. Other times writers may select the same category but with only partially overlap on the selected span (Table 7 Row 1 W3 vs W12). Looking at (Table 7 Row 4 and 5; W3 vs W12) there is a partial overlap in the selected span “and an unsettling sense of mystery that gnawed at me more than the inexplicable weight itself”. However, the selected categories are Purple Prose and Cliché respectively. Here again, it should be noted that Purple Prose is a style of writing that can be original or cliché, depending on its usage, context, and frequency. Not all elaborate writing is overused, but when certain ornate phrases or styles become too common, they can cross the line into cliché territory. W16 however did not edit this span.
@@ 6. AUTOMATIC DETECTION AND REWRITING OF LLM IDIOSYNCRASIES
Expert text editing can analyze and reduce LLM idiosyncrasies at a small scale, but automated methods are needed for resolution at a larger scale. Building on Hayes et al. (1987) and Scardamalia (1983), we develop techniques to separate detection and rewriting tasks, evaluating them using LAMP Corpus annotations. Given automated evaluation limitations for text editing(Dou et al., 2023), we conduct a large-scale preference annotation study with LAMP Corpus writers, comparing human and LLM-produced edits. To accommodate methods that require training samples, we split our data: 146 of 1057 LAMP Corpus paragraphs for training, the rest for testing.
@@ 6.1. Automatic detection of problematic spans in LLM-generated text
To evaluate various methods, we use the span-level precision metric, a common metric used in NLP tasks requiring comparison of extracted spans (Rajpurkar et al., 2016). Span-level precision measures the degree of overlap between predicted spans and reference or ground truth spans (in our case spans collected as a part of the editing process from writers). The overlap is measured at the character level, such that spans that partially overlap will get precision scores that reflect the amount of overlap between the two spans. High span-level precision indicates that the model is precise in identifying the correct boundaries of relevant text spans without over-predicting. We implement two precision metric variants: General and Categorical Precision. General Precision credits span selection regardless of category assignment, while Categorical Precision requires correct category assignment. We use a precision-based metric (like BLEU (Papineni et al., 2002)) rather than recall-based (like ROUGE (Lin, 2004)) as LLM-based methods tend to over-generate spans, which recall doesn’t penalize. Our focus is measuring overlap between generated and ground truth spans. Appendix A.2 provides a simplified example where General and Categorical Precision are computed, illustrating the suitability of these metrics for this evaluation setting.
General Precision Categorical Precision
@@
Detector LLM 2-shot 5-shot 25-shot 2-shot 5-shot 25-shot
@@
We implement few-shot LLM-based methods (Brown et al., 2020) that have demonstrated competitive performance on tasks across several disciplines, often using fewer than 100 examples  [^9] Model fine-tuning experiments on LAMP Corpus are left for future work. Our experiment varies few-shot examples (2, 5, and 25) with the 2-shot prompt in Appendix A.1 and tests Llama3.1-70B, GPT-4o, and Claude3.5-Sonnet. As part of the collection of the LAMP Corpus, 50 paragraphs were edited independently, and we computed General and Categorical Precision on this set to estimate expert-expert agreement. Table 9 summarizes results. The best General Precision (0.46) is achieved by Claude-3.5 Sonnet and GPT-4o with a 5-shot prompt, below the expert agreement level (0.57). LLM-based methods can identify problematic spans with significant expert overlap, but improvement is possible. Performance improves from 2-shot to 5-shot prompts but plateaus thereafter (Claude-3.5 Sonnet and GPT-4o achieve similar or better performance with 5-shot vs. 25-shot prompts).
@@
$[\textbf{\color[rgb]{1,0,0}tangible reality of a slower pace}_{\textsc{{\color[rgb]{0,0,1}Purple Prose)}}}]$ Here, in the $[\textbf{\color[rgb]{1,0,0}heart of progress}_{\textsc{{\color[rgb]{0,0,1}Cliche}}}]$, he was a mere observer,
detached from the legacy beneath the steel and concrete, his own inclinations buried beneath the layers of
@@ 6.3. Evaluating Automatic Editing of LLM-generated Text
To evaluate editing quality, we design an evaluation task where participants read three variants of a paragraph and rank them in terms of overall preference: (1) an unedited LLM-generated paragraph from the LAMP Corpus, (2) the Writer-edited version from the LAMP Corpus, and (3) an LLM-edited version using our pipeline to detect and rewrite problematic spans. We re-hired 12 of the 18 experts who had participated in creating the LAMP Corpus for this evaluation.
@@
Table 11 shows examples of LLM-edited paragraphs under both sub-conditions. In our pilot evaluation, we initially included all four conditions for annotation. However, ranking four paragraphs proved challenging for participants, especially when distinguishing between their second and third preferences. Based on this feedback, we redesigned the task to have participants judge only three conditions in each annotation. We always included the LLM-generated and Writer-edited paragraphs and alternated between including LLM-edited-oracle and LLM-edited-full paragraphs. We note that to obtain automatic edits of a paragraph, we used the same LLM that had originally been used to generate the paragraph. [^10] In other words, we used GPT-4o in the two-step pipeline to generate edits to paragraphs that were originally generated by GPT-4o. While not optimal, as a single LLM might offer slightly better detection and rewriting capabilities, this approach allows us to simplify the experiment conceptually and and also test our hypothesis if edits lead to overall better alignment without relying on a single model family. We assess if using an LLM in a multi-stage pipeline (drafting, problem detection, rewriting) can enhance overall writing quality. Future work could potentially optimize this editing pipeline further, possibly yielding better results for LLM-edited conditions.
@@
In summary, our experiment validates the potential benefit of automatic editing to improve writing quality: although automatic editing does match the quality of edits provided by professional writers, LLM-edited text is significantly preferred to LLM-generated text by expert writers (Design Principle 3). In other words, this experiment shows that LLMs can improve the quality of their writing in a fully automatic way, by first generating a draft, selecting problematic spans, and then rewriting such spans.
@@ 7.1. How is editing human writing different from LLM-generated text?
Editing human writing and LLM-generated text presents distinct challenges and requires different approaches. Human writing often contains nuanced expressions, personal style, and contextual references that reflect the author’s unique voice and experiences. Editors must preserve these elements while refining clarity, structure, and coherence. In contrast, LLM-generated text may lack consistent tone and exhibit repetitive patterns (Section 5). We asked writers to explain the differences in editing LLM-generated text compared to Human-written text. Several writers mentioned that LLM writing often required more extensive editing, mainly to remove unusual and sometimes nonsensical metaphors, inappropriate use of complex vocabulary that doesn’t fit the context and improving an overall tone that comes across as impersonal and mechanical. In a exchange W3 noted “I edit a lot of prose for my magazine but one thing that stuck with me as I was editing these paragraphs are the massive amount of cliché, histrionic descriptions, and direct exposition of intended meanings rather than effective representation. Indeed very strange”. They observed that the types of edits needed for LLM-generated text were often similar, but the sheer volume of necessary changes was higher than that with respect to human writing. LLM-generated content’s repetitive nature paradoxically made the human-editors feel robotic while trying to improve it.

### v5

@@ 4.4.7. Tense Inconsistency
Figure 5. (5(a)) the categories of edits they implement, and (5(b)) the relationship between writing quality scores and error categories. (5(c)) Distribution of semantic similarity scores for the edits in the dataset
@@ 4.5. Collecting Edits on LLM Generated responses
With the finalized taxonomy of edits, we next conducted a larger-scale annotation study. The purpose of this study was to collect edits from writers on LLM-generated responses, categorizing them according to the established taxonomy. We validated the taxonomy’s comprehensiveness by consulting writers who were not part of the formative study but participated in the editing task. These writers had the option to select an ”Other” category and provide its name if an edit didn’t fall into any of the seven established categories. The writers rarely chose the Other category, doing so in only 10 out of 8,035 cases. These rare exceptions fell into categories such as “Repetitive Sentence Structure,” “Confusing, Unclear or Incomplete Action/Meaning,” and “Mixed Metaphors”. We also separately asked the writers if the categories encompassed all traits they encountered while editing these paragraphs and whether they would suggest any additions. Through email exchanges, all writers confirmed the taxonomy’s comprehensiveness.
@@
After completing their edits, participants assigned two scores to the sample: an Initial Writing Quality Score (IWQS) for the original response quality, and a Final Writing Quality Score (FWQS) for the post-edit quality. Both used a 1-10 scale, with 1 being the lowest and 10 being the highest quality. The scores were incorporated to add a quantitative dimension to the qualitative process of editing. Additionally, the self-reported writing quality scoring system serves as a signal for writers to recognize their own improvements, set personal goals, and develop intrinsic motivation for enhancing their work.
@@ 5.1. Overall Statistics
We analyze the editing process by examining edit operations: insertion, deletion, or replacement. An edit is an insertion if it deletes no characters or adds 40+ characters net. Conversely, it’s a deletion if it adds no characters or removes 40+ characters net. All other edits are tagged as replacements. We choose a threshold of 40 characters (roughly 10 words), to avoid labeling edits with minor length changes (“a” to “the”) as insertions or deletions. Figure 4(a) shows edit operations by participant for each paragraph. Replacements are most frequent (74%), followed by deletions (18%) and insertions (8%). Editing styles vary: some participants primarily use replacements (W2, W9, W10, W16, W18), while others employ deletions more often (W1, W5, W7, W8, W13, W17). Insertions are uncommon across all participants. To quantify meaning-preserving vs. meaning-changing edits, we calculate semantic similarity between original and edited text using BERT score(Zhang et al., 2019). Using a threshold of 0.6  [^7] This threshold was decided by manually analyzing 100 edits, we classify edits with similarity ¿ 0.6 as meaning-preserving. Of 6468 non-deletion edits, 70% are meaning-preserving, with the rest meaning-changing. This finding supports our Design Principle 2. Figure 5(c) shows the distribution of the semantic similarity scores for the edits in LAMP.  [^8] However we note that with a threshold of 0.6, this appears to be quite a fuzzy/arbitrary distinction rather than a clear separation. While the distribution is bimodal (with peaks around 0.6 and 0.75), these modes don’t align well with the 0.6 threshold. This suggests that meaning preservation versus meaning change exists more on a continuum rather than as two clearly distinct categories.
@@ 5.2. Writers differ greatly in the amount of editing they do: But to what extent?
Table 7 shows how sometimes writers select the exact same problematic span but assign different categories. A span that both W3 and W16 selected “, the numbers glaring back at me like an unsolvable riddle” (Table 7 Row 1) was categorized differently (Table 7 Row 2). Yet both categorizations can be correct interpretations. When one relies on overused phrases or clichés, they often state the obvious or provide information that readers can easily infer implicitly. This results in redundant or superfluous exposition that doesn’t add value to the narrative. Other times writers may select the same category but with only partial overlap on the selected span (Table 7 Row 1 W3 vs W12). Looking at (Table 7 Row 4 and 5; W3 vs W12) there is a partial overlap in the selected span “and an unsettling sense of mystery that gnawed at me more than the inexplicable weight itself”. However, the selected categories are Purple Prose and Cliché respectively. Here again, it should be noted that Purple Prose is a style of writing that can be original or cliché, depending on its usage, context, and frequency. Not all elaborate writing is overused, but when certain ornate phrases or styles become too common, they can cross the line into cliché territory. W16 however did not edit this span.
@@ 6. AUTOMATIC DETECTION AND REWRITING OF LLM IDIOSYNCRASIES
While several automated editing approaches exist for improving LLM outputs at scale, we evaluate their effectiveness specifically for addressing idiosyncrasies in creative writing. Building on Hayes et al. (1987) and Scardamalia (1983), we develop techniques to separate detection and rewriting tasks, evaluating them using LAMP Corpus annotations. Given automated evaluation limitations for text editing(Dou et al., 2023), we conduct a large-scale preference annotation study with LAMP Corpus writers, comparing human and LLM-produced edits. To accommodate methods that require training samples, we split our data: 146 of 1057 LAMP Corpus paragraphs for training, the rest for testing.
@@ 6.1. Automatic detection of problematic spans in LLM-generated text
To evaluate various methods, we use the span-level precision metric, a common metric used in NLP tasks requiring comparison of extracted spans (Rajpurkar et al., 2016). Span-level precision measures the degree of overlap between predicted spans and reference or ground truth spans (in our case spans collected as a part of the editing process from writers). The overlap is measured at the character level, such that spans that partially overlap will get precision scores that reflect the amount of overlap between the two spans. High span-level precision indicates that the model is precise in identifying the correct boundaries of relevant text spans without over-predicting. We implement two precision metric variants: General and Categorical Precision. General Precision credits span selection regardless of category assignment, while Categorical Precision requires correct category assignment. We use a precision-based metric (like BLEU (Papineni et al., 2002)) rather than recall-based (like ROUGE (Lin, 2004)) as LLM-based methods tend to over-generate spans, which recall doesn’t penalize. Our focus is measuring the overlap between generated and ground truth spans. Appendix A.2 provides a simplified example where General and Categorical Precision are computed, illustrating the suitability of these metrics for this evaluation setting.
General Categorical
@@
Detector LLM n=2 n=5 n=25 n=2 n=5 n=25
@@
$\textbf{\color[rgb]{1,0,0}tangible reality of a slower pace}_{\textsc{{\color[rgb]{0,0,1}Purple Prose)}}}]$ Here, in the $[\textbf{\color[rgb]{1,0,0}heart of progress}_{\textsc{{\color[rgb]{0,0,1}Cliche}}}]$, he was a mere observer,
[[[[detached from the legacy beneath the steel and concrete, his own inclinations buried beneath the layers of
@@
We implement few-shot LLM-based methods (Brown et al., 2020) that have demonstrated competitive performance on tasks across several disciplines, often using fewer than 100 examples  [^10] Model fine-tuning experiments on LAMP Corpus are left for future work. Our experiment varies the number of few-shot examples (2, 5, and 25) with the 2-shot prompt in Appendix A.1 and tests Llama3.1-70B, GPT-4o, and Claude3.5-Sonnet. As part of the collection of the LAMP Corpus, 50 paragraphs were edited independently, and we computed General and Categorical Precision on this set to estimate expert-expert agreement. Table 9 summarizes results. The best General Precision (0.46) is achieved by Claude-3.5 Sonnet and GPT-4o with a 5-shot prompt, below the expert agreement level (0.57). LLM-based methods can identify problematic spans with significant expert overlap, but improvement is possible. Performance improves from 2-shot to 5-shot prompts but plateaus thereafter (Claude-3.5 Sonnet and GPT-4o achieve similar or better performance with 5-shot vs. 25-shot prompts).
@@ 6.3. Evaluating Automatic Editing of LLM-generated Text
To evaluate editing quality, we design an evaluation task where participants read three variants of a paragraph and rank them in terms of overall preference: (1) an unedited LLM-generated paragraph from the LAMP Corpus, (2) the Writer-edited version from the LAMP Corpus, and (3) an LLM-edited version using our pipeline to detect and rewrite problematic spans. 6 out of 18 of our experts were unavailable during the preference annotations. Some of them were busy with their full-time jobs while others wanted to take a break after the edit task. Given the concern for proper turn around we re-hired 12 of the 18 experts who had participated in creating the LAMP Corpus for this evaluation.
@@
Table 11 shows examples of LLM-edited paragraphs under both sub-conditions. Although the examples here showcase ‘Replacement’ and ‘Deletion’ edits, analysis on a larger corpus of 200 LLM-edited paragraphs confirms that the automatic editing pipeline mirrors the edit type distribution of expert editors (shown in Figure 4(a)): roughly 65% ‘replace‘ edits, 25% ‘delete‘ edits, and 10% ‘insert‘ edits. This surface-level analysis confirms the automatic pipeline’s ability to mimic the editing process of experts measured in terms of relative proportions of character insertion and deletion.
In our pilot evaluation, we initially included all four conditions for annotation. However, ranking four paragraphs proved challenging for participants, especially when distinguishing between their second and third preferences. Based on this feedback, we redesigned the task to have participants judge only three conditions in each annotation. We always included the LLM-generated and Writer-edited paragraphs and alternated between including LLM-edited-oracle and LLM-edited-full paragraphs. While traditionally pairwise rankings tend to be easier for humans to annotate, three-way ranking allowed direct comparison between the key conditions we wanted to evaluate (LLM-generated v.s. Writer-edited v.s. LLM-edited), making the relative preferences clearer in a single evaluation rather than having to combine multiple pairwise comparisons. Additionally, we want to avoid any familiarity effects or biases that could arise from evaluators reading the same text multiple times, which would likely occur when collecting overlapping pairwise comparisons.
We note that to obtain automatic edits of a paragraph, we used the same LLM that had originally been used to generate the paragraph. [^11] In other words, we used GPT-4o in the two-step pipeline to generate edits to paragraphs that were originally generated by GPT-4o. While not optimal, as a single LLM might offer slightly better detection and rewriting capabilities, this approach allows us to simplify the experiment conceptually and also test our hypothesis if edits lead to overall better alignment without relying on a single model family. We assess if using an LLM in a multi-stage pipeline (drafting, problem detection, rewriting) can enhance overall writing quality. Future work could potentially optimize this editing pipeline further, possibly yielding better results for LLM-edited conditions.
@@
In summary, our experiment validates the potential benefit of automatic editing to improve writing quality: although automatic editing does match the quality of edits provided by professional writers, LLM-edited text is significantly preferred to LLM-generated text by expert writers (Design Principle 3). In other words, this experiment shows that LLMs can improve the quality of their writing in a fully automatic way, by first generating a draft, selecting problematic spans, and then rewriting such spans. These results align with previous findings (Wadhwa et al., 2024; Pan et al., 2024b; Gao et al., 2024b; Dwivedi-Yu et al., 2022) showing that iterative refinement and editing can improve LLM outputs.
Genre Category Original Span LLM edited span
Literary
Fiction
Lack of Specificity and Detail
seems to embody the city’s frustrations
and disappointments
seems to embody the city’s frustrations
and disappointments, his eyes a deep
well of anger and desperation, his face a
topographic map of the city’s corruption
and neglect.
Food
Writing
Gone are the individual ramekins,
replaced by a single, generous vessel
that invites sharing.
Gone are the individual ramekins, replaced
by a deep-dish pie plate that invites
communal indulgence. The golden crust
cradles a sea of silky custard, its surface
a glassy expanse of caramelized sugar.
Literary
Fiction
Purple Prose
confusion clouded his understanding
like the haze hanging over the park
he remained confused
Food
Writing
It’s the democratization of ingredients
that sets the bánh mì apart
It’s the variety of ingredients that
sets the bánh mì apart
Internet
Advice
But pain has a way of cracking us open,
making space for new possibilities we
couldn’t have imagined before.
But pain can open us up to new
possibilities.
Travel
Writing
Poor Sentence Structure
He tips generously but modestly, as if
acknowledging a pact of respect
He tips generously yet with restraint,
honoring a mutual respect
Literary
Fiction
The gulf between her vision for her
future and her parents’ hopes wasn’t
something that could be bridged with
words alone, and Bella realized that
sometimes growth requires difficult
choices.
Bella realized that her vision for the
future and her parents’ hopes were
irreconcilable. Growth, she understood,
often demands difficult choices.
Internet
Advice
Unnecessary Exposition , a silent reminder of everything it once meant -
Personal
Essay
Now, when life deals its inevitable blows, I
think of those stones. I remember that
sometimes it’s the cracks that let the light in,
and that even the most jagged edges can be
smoothed by time and patience.
Now, when life deals its inevitable
blows, I think of those stones.
Travel
Writing
Cliché
As the weekend comes to a close, reflect on the
kaleidoscope of experiences you’ve had, and
how each neighborhood has contributed to the
vibrant tapestry that is Los Angeles.
As the weekend comes to a close, you’ll
have a new appreciation for the city’s
diversity.
Food
Writing
a dish that feels both familiar and unexpected,
much like discovering a hidden alley in a
well-known city. It’s
a familiar dish with a surprising
twist. It’s
Literary Fiction , watching his chest rise and fall with labored breaths -
I noticed a vase of wilting flowers on the nightstand,
forgotten in the weight of more pressing concerns
The flowers on the nightstand had wilted
Table 13. LLM edited spans from different categories across different sub-genres of writing
6.4. Qualitative Insights into LLM edits
Given how LLM-edited paragraphs are often preferred over default LLM-generated ones and sometimes even Writer edited paragraphs, we inspect if there are specific categories of edits where LLMs perform better, more so if that caters to a certain genre. Table 13 shows examples of original spans edited by LLMs that vary across genre and category. In general, there isn’t a domain or category where LLM edits are better than the rest except for Purple Prose. To edit Purple Prose models mostly need to learn how to simplify the text (i.e., generate a paraphrase in simple language). LLMs are effective at Sentence Simplification (Dou et al., 2024) so this doesn’t come across as a surprising finding. Categories like Poor Sentence Structure and Unnecessary Exposition deliver mixed results. Sometimes they are effective. For instance Row 7 in Table 13 shows how the model improves the structure by breaking the run-on very long sentence. However, for the other example in Row 6, the human edit is Then he hands his card, always leaving a 22% tip while the model edit is a mere paraphrase. Similarly, under the Unnecessary Exposition category, the LLM is very good at editing sentences with the following structure [Main clause] [comma] [exposition] where they simply remove the exposition (as learned from Writer edits). However, the bigger issue in this category is revealing the subtext. When a given input span doesn’t conform to the structure above or consists of multiple sentences (Row 9, Table 13) LLMs don’t align with human edits. For Cliché, models mostly rewrite it in simpler language typically with fewer words which while being effective isn’t what humans do. This also shows the fundamental overlap between Cliché and Purple Prose as shown in Section 5.2. Finally, models are mostly ineffective at addressing Lack of Specificity and Detail (more details in section 7.2). These findings demonstrate the challenge of editing text across any domains that require deeper emotional resonance or cultural commentary and the importance of rethinking the design of alignment for tasks with subjective or fuzzy rewards.
@@ 7.1. How is editing human writing different from LLM-generated text?
Editing human writing and LLM-generated text presents distinct challenges and requires different approaches. Human writing often contains nuanced expressions, personal style, and contextual references that reflect the author’s unique voice and experiences. Editors must preserve these elements while refining clarity, structure, and coherence. In contrast, LLM-generated text may lack consistent tone and exhibit repetitive patterns (Section 5). We asked writers to explain the differences in editing LLM-generated text compared to Human-written text. Several writers mentioned that LLM writing often required more extensive editing, mainly to remove unusual and sometimes nonsensical metaphors, inappropriate use of complex vocabulary that doesn’t fit the context, and improving an overall tone that comes across as impersonal and mechanical. In an exchange, W3 noted “I edit a lot of prose for my magazine but one thing that stuck with me as I was editing these paragraphs are the massive amount of cliché, histrionic descriptions, and direct exposition of intended meanings rather than effective representation. Indeed very strange”. They observed that the types of edits needed for LLM-generated text were often similar, but the sheer volume of necessary changes was higher than that of human writing. LLM-generated content’s repetitive nature paradoxically made the human editors feel robotic while trying to improve it. By the edits that catered to deep emotional resonance or cultural commentary,